OpenStudy (anonymous):

Please help me solve for x , and indentify if its extraneous. square root only over over x-95 ,+9 = 6

OpenStudy (whpalmer4):

\[\sqrt{x-95}+9=6\]?

OpenStudy (anonymous):

Yes

OpenStudy (whpalmer4):

You're sure about that?

OpenStudy (anonymous):

yes because my teacher is asking me to substitute my birth date and year into this equation

OpenStudy (anonymous):

\[\sqrt{x-95}= -3\]

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

oh okay. so that's what I start with?

OpenStudy (whpalmer4):

Problem is that there's no real solution… \[\sqrt{x-95}+9=6\]\[\sqrt{x-95}=-3\]There's no real number that satisfies that...

OpenStudy (anonymous):

x-95=9 and yess

OpenStudy (anonymous):

now just add 95 on both sides

OpenStudy (anonymous):

okay so maybe ill just tell her to screw,and I cant do it.yay.

OpenStudy (whpalmer4):

Maybe that's what your instructor hopes you'll find...

OpenStudy (anonymous):

ok hold on, let me attempt this

OpenStudy (anonymous):

you get x=92

OpenStudy (anonymous):

I got x = 104?

OpenStudy (anonymous):

ooops yes you are correct

OpenStudy (anonymous):

hallelujah

OpenStudy (anonymous):

i was looking at the previous forms of this equation

OpenStudy (anonymous):

thank you gentlemen for your help

OpenStudy (anonymous):

nice! no problem

OpenStudy (whpalmer4):

you aren't done yet, though — you need to test out your solution to see if it works.

OpenStudy (anonymous):

oh no

OpenStudy (whpalmer4):

\[\sqrt{104-95} +9 = 6\]Can you simplify that?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

@whpalmer4 thats right, forgot about that. good thing we are both here

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok I got 9=9 at the end

OpenStudy (anonymous):

so I hope im right

OpenStudy (whpalmer4):

no, you don't get that… what is the value of the square root?

OpenStudy (anonymous):

uhm, im not sure

OpenStudy (whpalmer4):

then how could you know that you got the right thing?

OpenStudy (whpalmer4):

\[\sqrt{104-95} = \]

OpenStudy (anonymous):

u get \[\sqrt{9}=6\]

OpenStudy (anonymous):

or just 3

OpenStudy (anonymous):

okay, so I minused 104 from 95, got 9 plus 9 = 6 then I squared both and took 9 out of the square root

OpenStudy (anonymous):

then I did 6-9 which is -3 but squared is 9 so I got 9=9

OpenStudy (anonymous):

ohhhh no shes right

OpenStudy (anonymous):

it can be plus or minus 3

OpenStudy (anonymous):

or realy

OpenStudy (anonymous):

+3+9=6 is wrong but -3+9=6

OpenStudy (anonymous):

* oh really

OpenStudy (anonymous):

ok well good deal

OpenStudy (whpalmer4):

no, without a sign in front of the radical, it means the principal root, which is the positive one.

OpenStudy (whpalmer4):

leaving us with \[3+9=6\]

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

INSTEAD OF SQUARING BOTH SIDES SHE CAN SIMPLIFY THE SQUARE ROOT OF 9 with = plus or minus 3

OpenStudy (anonymous):

*which

OpenStudy (whpalmer4):

so x=104 is an extraneous solution. we introduced it when we squared both sides of the equation.

OpenStudy (whpalmer4):

only \[-\sqrt{x-95}+9=6\]has a solution

OpenStudy (anonymous):

I thought it is only extraneous if theres a negative sign in front.

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

are you sure @whpalmer4

OpenStudy (anonymous):

okay

OpenStudy (whpalmer4):

110% sure.

OpenStudy (whpalmer4):

no, extraneous means "it doesn't work" :-)

OpenStudy (anonymous):

im already lost at 1:43

OpenStudy (anonymous):

I don't understand why theres a -24x

OpenStudy (whpalmer4):

\[(2x-6)(2x-6) = 2x*2x-6*2x-6*2x+36 = 4x^2-24x+36\]